OpenStudy (anonymous):

someone prove this- if we have two sequences s(n) and l(n) with limits 'a' and 'b' then limit s(1).l(n)+s(2).l(n-1)+s(3).l(n-2)+_ _ _ _ +s(n).l(1)/n as n tends to infinity is equal to a.b

OpenStudy (anonymous):

the one who solves this problem will be appreciated with medal

OpenStudy (anonymous):

someone help me pppppplllllllllllzzzzzzzzzz

OpenStudy (blockcolder):

\[\lim_{n\rightarrow\infty} s(n)=a; \lim_{n\rightarrow\infty} l(n)=b\] To prove: \[\lim_{n\rightarrow\infty} \frac{s(1)l(n)+s(2)l(n-1)+\cdots +s(n)l(1)}{n}=ab\] Is this the statement of your problem?

OpenStudy (anonymous):

yes if you could please help me in proving this statement